OpenStudy (anonymous):

Which is the solution of the equation below, rounded to the nearest tenth? 1.6x – 3.38 = 5.81x + 4.67

OpenStudy (anonymous):

answers could be –0.2 –0.3 –1.9 –2.0

OpenStudy (anonymous):

3rd one

OpenStudy (anonymous):

-1.9

jigglypuff314 (jigglypuff314):

solve for x first so 1.6x – 3.38 = 5.81x + 4.67 you would subtract 4.67 from both sides - 4.67 - 4.67 1.6x - 8.05 = 5.81x + 0 then subtract 1.6x on both sides -1.6x -1.60x 0 - 8.05 = 4.21x then divide both sides by 4.21 /4.21 /4.21 - 1.912114 = x then round to the tenth place so -1.9 = x

OpenStudy (anonymous):

told u i was correct retricewell as jigglypuff314